# Partial identification and unmeasured confounding with multiple treatments and multiple outcomes

### v2

@@ Appendix C Estimation and inference on bounds
In this section, we describe our estimation procedure for the models (1)–(3), as well as an inferential procedure that accounts for sampling uncertainty when working with partial identification regions or bounds. For now, we assume $\Sigma_{u}=I_{m}$, $\Sigma_{t|u}=\sigma^{2}_{t|u}I_{k}$, and $\Sigma_{y|t,u}=\sigma^{2}_{y|t,u}I_{q}$, though it is relatively straightforward to adapt these strategies to non-equal variances.
@@ C.1 Estimation of factor model parameters
The parameters corresponding to the treatment model, $\mu_{u|t}$ and $\Sigma_{u|t}$, can be estimated using factor analysis on the observed treatment matrix. From (2), we have that
@@
The outcome model follows in a very similar way, though it requires an additional step where we first remove the effect of the treatment on the outcome. Specifically, we first estimate $\check{g}(\cdot)$ using a regression of $Y$ on $T$, and compute $Y-\check{g}(T)$. We then perform the same factor analysis steps as for the treatment model, but on the correlation matrix of $Y-\check{g}(T)$, which will provide an estimate $\widehat{\Gamma}$ of the unknown factor loadings. Once we have obtained estimates $\widehat{B}$ and $\widehat{\Gamma}$, we immediately obtain estimates of any identified quantity such as the partial $R^{2}$ values between the unmeasured confounders and treatment or outcome, as well as the bias bounds derived in (9).
One key component of factor analysis is determining how many factors to retain, as it is generally unknown in real applications. This is a well studied problem in the factor analysis literature, and there are a number of existing methods for determining the number of factors. Possibilities include the very simple structure criterion with complexity 1 and 2 (Revelle & Rocklin 1979), the minimum average partial criterion (Velicer 1976), BIC, adjusted BIC, including the number of factors with eigenvalues greater than those of random data (parallel analysis), or the number of eigenvalues greater than 1. We leave discussion to the merits of each of these approaches to the existing literature on factor analysis, though empirically we have found both BIC and factor analysis to work reasonably well in the simulations explored in our work. We recommend favoring too many factors instead of too few factors as this will generally lead to more conservative inferences. In the analysis of PM2.5 components, we used the largest number of factors allowable with 6 exposures and 6 outcomes, which is $m=3$.

### v3

@@ Appendix C Estimation and inference on bounds
In this section, we describe our estimation procedure for the required statistical models, as well as an inferential procedure that accounts for sampling uncertainty when working with partial identification regions or bounds. For now, we assume $\Sigma_{u}=I_{m}$, $\Sigma_{t|u}=\sigma^{2}_{t|u}I_{k}$, and $\Sigma_{y|t,u}=\sigma^{2}_{y|t,u}I_{q}$, though it is relatively straightforward to adapt these strategies to non-equal variances.
@@ C.1 Estimation of factor model parameters
The parameters corresponding to the treatment model, $\mu_{u|t}$ and $\Sigma_{u|t}$, can be estimated using factor analysis on the observed treatment matrix. From the treatment model, we have that
@@
The outcome model follows in a very similar way, though it requires an additional step where we first remove the effect of the treatment on the outcome. Specifically, we first estimate $\check{g}(\cdot)$ using a regression of $Y$ on $T$, and compute $Y-\check{g}(T)$. We then perform the same factor analysis steps as for the treatment model, but on the correlation matrix of $Y-\check{g}(T)$, which will provide an estimate $\widehat{\Gamma}$ of the unknown factor loadings. Once we have obtained estimates $\widehat{B}$ and $\widehat{\Gamma}$, we immediately obtain estimates of any identified quantity such as the partial $R^{2}$ values between the unmeasured confounders and treatment or outcome, as well as the bias bounds seen in Theorem 1.
One key component of factor analysis is determining how many factors to retain, as it is generally unknown in real applications. This is a well studied problem in the factor analysis literature, and there are a number of existing methods for determining the number of factors. Possibilities include the very simple structure criterion with complexity 1 and 2 (Revelle and Rocklin,, 1979), the minimum average partial criterion (Velicer,, 1976), BIC, adjusted BIC, including the number of factors with eigenvalues greater than those of random data (parallel analysis), or the number of eigenvalues greater than 1. We leave discussion to the merits of each of these approaches to the existing literature on factor analysis, though empirically we have found both BIC and factor analysis to work reasonably well in the simulations explored in our work. We recommend favoring too many factors instead of too few factors as this will generally lead to more conservative inferences. In the analysis of PM2.5 components, we used the largest number of factors allowable with 6 exposures and 6 outcomes, which is $m=3$.